Find the vertex: f(x)= -(x+6)^2-8
Solve: 5x^2=10
Find the vertex: f(x) = 5(x+5)^2+2
Find the vertex: f(x) = 3x^2-18x+8

Answer by stanbon(75887)

Find the vertex: f(x)= -(x+6)^2-8
Ans: (-6,-8)
---------------------------
Solve: 5x^2=10
x^2= 2
x = sqrt(2) or x = -sqrt(2)
--------------------------------
Find the vertex: f(x) = 5(x+5)^2+2
Ans: (-5,2)
-----------------------------------
Find the vertex: f(x) = 3x^2-18x+8
Vertex occurs where x = -b/(2a) = 18/(6) = 3
f(3) = 3*9-18*3+8 = 27-54+8 = 35-54 = -21
Ans: (3,-21)
=================
Cheers,
Stan H.
=================